The problems listed below were sent a few min. ago, but I am not sure they went through. I am helping my grandson.. but I need help. thank you very much
x/9=15/2 9.47/p=3.33/1 20/8.7=5.8/m

Answer by checkley71(8403)

X/9=15/2 CROSS MULTIPLE GIVES 2X=135 THEN X=135/2 OR X=67.5
9.47/P=3.33/1 CROSS MULTIPLY GIVES 3.33P=9.47 THEN P=9.47/3.33 OR P=2.84
20/8=5.8/M CROSS MULTIPLY GIVES 20M=46.4 OR M=46.4/20 OR M=2.32
HOPE THIS HELPS GRAMPS.